The table below shows the number of male and female students enrolled in nursing at a university for a certain semester. A student is selected at random. Complete parts (a) through (d). 


Nursing majorsNon-nursing majorsTotal
Males9810161114
Females60017292329
Total69627453443

(c) Find the probability that the student is not female or a nursing major. 

P(not being female or being a nursing major)=_______ 

Round to the nearest thousandth as needed.) 

(d) Are the events "being male" and "being a nursing major" mutually exclusive? Explain. 

A. Yes, because one can't be male and a nursing major at the same time. 

B. Yes, because there are 98 males majoring in nursing 

C. No, because there are 98 males majoring in nursing 

D. No, because one cant be male and a nursing mayor at the same time.
Answer #2

SOLUTION :


c.


P( Not female or Nursing majors)

= P(Male or Nursing majors) 

= P(Male) + P(nursing majors) - P(both)

= 1114/3443 + 698/3443 - 98/3443

= (1114 + 698 -  98) / 3443

= 1714/3443

=  0.498 (ANSWER).



d.


Option C : No, since there are 98 males in Nursing majors.(ANSWER).






answered by: Tulsiram Garg
Answer #1

P(not female or nursing major ) = P(Not female) + P(nursing major) - P( not female and nursing major)

= 1114/3443 + 698/3443 - 98/3443

=1714/3443

=0.4978

P(not female or nursing major ) = 0.498

d)

No, because there are 98 males majoring in nursing